What is the y-intercept of y = 1/2x + b if it passes through (-8,-4)?
a. m=0
b. b=0
c. b=-4
d. Insufficient information given

Answer by oscargut(2103)

y = 1/2x + b (y-intercept=f(0)=b)
it passes through (-8,-4) so -4=1/2(-8)+b then -4=-4+b then b=0
so option b is the correct